Real and nominal rates interest Zane Perelli currently has $116 that he can spend today on socks costing $2.90 each. Alternatively, he could invest the $116 in a risk-free U.S. Treasury security that is expected to earn a 12% nominal rate of interest. The consensus forecast of leading economists is a 5% rate of inflation over the coming year. a. How many socks can Zane purchase today? b. How much money will Zane have at the end of 1 year if he forgoes purchasing the socks today and invests his money instead? (Ignore taxes.) c. How much would you expect the socks to cost at the end of 1 year in light of the expected inflation? d. Use your findings in parts b and c to determine how many socks (fractions are OK) Zane can purchase at the end of 1 year. In percentage terms, how many more or fewer socks can Zane buy at the end of 1 year? e. What is Zanes real rate of return over the year? How is it related to the percentage change in Zanes buying power found in part d? Explain.
Answer #1
Zane's current available money = $116
Total number Socks he can purchase = Total available money/Price of one socks
Price of one Socks = $2.90
a. Total number Socks Zane can purchase = $116/$2.90 = 40 Socks
Expected Nominal Rate of return upon investment = 12%
Inflation rate = 5%
Real Rate of Interest = 1+ Nominal rate/1+ Inflation rate - 1
= (1+0.12/1+0.05)-1
= 6.666% or 6.67%
b. If Zane invests in Risk free U.S. Treasury, he earns 6.67% (Real) rate of Interest on $116, which is
=$(116+0.667*116) = $193.4
c. Cost of 40 Socks at the end of year with 5% inflation rate
= $(2.90+2.90*.05)*40
= $121.8 for 40 Socks
d. Considering the Year end price of One Socks, which is $(2.90+*.05*2.90) = $3.045
Zane can buy $193.4/$3.045 Socks at the end of the year
= 63.52 Socks
So, Zane can buy (63.52-40) more Socks
= 23.52 Socks
or 23.52/40*100 = 58.8% more
Zane's Real Rate of Return is 6.67%, which co-relates by 58.8% increase in the Purchasing Power of Zane overall a period of 1 year